OpenStudy (anonymous):

find the requested function value. f(x)=3x+7, g(x)=-7x^2-4x-2 Find (fxg)(3)

OpenStudy (anonymous):

find \(f(3)\), then find \(g(3)\) then multiply the results is it clear how to find these?

OpenStudy (anonymous):

i will do the hard one \[g(3)=-7\times 3^2-4\times 3-2=-7\times 9-12-2=-77\]

OpenStudy (anonymous):

@satellite73 yeah man thats what I've been trying to do, but the only options are a) -178 b)-224 c)-98 d)-1858